# RealmDreamer: Text-Driven 3D Scene Generation with Inpainting and Depth Diffusion

### v1

@@ Abstract
We introduce RealmDreamer, a technique for generation of general forward-facing 3D scenes from text descriptions. Our technique optimizes a 3D Gaussian Splatting representation to match complex text prompts. We initialize these splats by utilizing the state-of-the-art text-to-image generators, lifting their samples into 3D, and computing the occlusion volume. We then optimize this representation across multiple views as a 3D inpainting task with image-conditional diffusion models. To learn correct geometric structure, we incorporate a depth diffusion model by conditioning on the samples from the inpainting model, giving rich geometric structure. Finally, we finetune the model using sharpened samples from image generators. Notably, our technique does not require video or multi-view data and can synthesize a variety of high-quality 3D scenes in different styles, consisting of multiple objects. Its generality additionally allows 3D synthesis from a single image.
Figure 1: A scene created by our technique on the left with prompt above compared to baseline ProlificDreamer [66] on the right. RealmDreamer generates 3D scenes from text prompts such as the one above. Using pretrained 2D inpainting and depth diffusion models, our method achieves state-of-the-art results with parallax, detailed appearance, and high-fidelity geometry.
@@ 1 Introduction
Text-based 3D scene synthesis has the potential to revolutionize 3D content creation by democratizing synthesis of high-fidelity 3D environments. With the increasing popularity of virtual reality and other spatial computing interfaces, this task has recently become even more important. Creating novice-friendly systems is made difficult by the long iteration times required for existing text-to-3D methods and restriction to simple object-level data or panoramas. An ideal 3D generation tool should offer users ownership over more aspects of the generation process, and fundamentally, the ability to create 3D scenes for immersive content creation and experiences (see Fig. 1).
The scarcity of scene-level 3D data makes generating 3D representations highly challenging. With virtually no scene-level 3D priors available, inconsistent 2D priors, such as CLIP [47] and Stable Diffusion [51], become crucial. Consequently, creating consistent gradients across multiple views for cohesive scene generation remains significantly difficult.
Trained on a large corpus of data, text-conditioned 2D diffusion models have shown excellent use as general purpose priors for a variety of tasks in computer vision, such as generation [72], editing [38, 43], classification [13], and segmentation [68]. Despite being a 2D prior over images, these models have also found application in 3D settings, such as enabling view synthesis [15, 75, 15] or editing [19] in the absence of any 3D training data. Most relevant to our task, some methods create 3D representations from text [46, 66] by optimizing the diffusion prior with respect to a 3D representation. However, due to the high complexity of 3D scenes, these methods have thus far been restricted to object-level data or panoramas with incorrect learned geometry (see Fig. 2).
To address this problem, we introduce RealmDreamer, a method for high-fidelity generation of 3D scenes from text prompts. Our key insight is to use pretrained inpainting and depth priors with a robust initialization of a 3D Gaussian Splatting model [30], to produce scenes that can be rendered from a wide baseline. Concretely, our contributions are the following:
@@
    A robust scene initialization for scene-level 3DGS with 2D priors (Sec. 4.1)
@@
    A framework for learning consistent 3D with 2D inpainting priors, within an occluded volume (Sec. 4.2)
@@
    A method for distillation from diffusion-based depth estimators (Sec. 4.3)
@@
    State-of-the-art results in text-based generation of 3D scenes (see Fig. 7)
Figure 2: Our method, compared to the state-of-the-art ProlificDreamer [66], shows significant improvements. ProlificDreamer’s public results, with outward-looking cameras from a scene-independent sphere, result in poor depth quality. Even with a complex training trajectory, ProlificDreamer yields unsatisfactory geometry and oversaturated renders. In contrast, our approach produces visually appealing images with realistic geometry.
@@ 2 Related Work
Text-to-3D. The first methods for text to 3D generation were based on retrieval from large databases of 3D assets [14, 8, 7]. Subsequently, learning-based methods have dominated. Using GANs, Text2Shape [9] learned to create textured 3D representations from text descriptions with GANs [1]. However, due to the dearth of diverse paired text and 3D data, more recent methods leverage 2D priors. Methods like DreamFields [25] and CLIP-Forge [54] leveraged CLIP [47] priors for text-to-3D. The most successful methods, however, have recently relied on the large-scale priors learned from text-to-image diffusion models [51, 53]. The seminal DreamFusion [46] and SJC [65] proposed to distill knowledge from such models into a 3D representation through score distillation sampling (SDS). There have been many follow-up works on this  [10, 32, 39, 36], such as ProlificDreamer [66] which improves SDS by training an additional view-dependent Low-Rank Adaptation [24] for better 3D consistency.
View synthesis with diffusion and 3D inpainting. Since the introduction of SDS, many papers have leveraged the capabilities of 2D diffusion models for 3D tasks. View synthesis from a single image may be accomplished with similar losses to SDS [75, 15]. 3DiM [67] learns to directly sample views conditioned on relative camera poses and an input image. Recent methods have pushed this direction further by increasing dataset size and exploring better conditioning architectures [34, 55, 56, 35]. Other methods perform diffusion in 3D space [44] or in a 3D latent space [26]. Finally, some methods condition denoising directly on renderings from 3D consistent models [17, 6] for inpainting in a multi-view consistent manner. However, all of the methods mentioned rely on object-level data, making scene-level synthesis infeasible. Separately, some methods focus on inpainting in 3D [42, 41], yet their results are limited to small baselines and inpainting masks. iNVS [27] also leverages inpainting diffusion models for view synthesis of objects from a single image.
Concurrent work. There are numerous concurrent works in the fast-moving text-to-3D space. Due to their number, we note only the most relevant concurrent works and clarify our differences from them. Some recent methods [70, 11, 63, 31] propose to utilize 3DGS representations as a backbone for text-to-3D, focusing on object-centric data. Text2NeRF [71] optimizes a NeRF for scene-level view synthesis, but its results are limited to narrow baselines with little parallax. Scenewiz3D [73] improves ProlificDreamer with the incorporation of RGBD priors, but the results lack realism. LucidDreamer [12] additionally attempts to tackle the task of 3D scene generation, but its results are mostly limited to simple camera trajectories such as those exhibiting strictly rotation or translation. In contrast, our method leverages inpainting and depth priors for scene synthesis with non-trivial camera trajectories.
@@ 3 Preliminaries
The splats $\{\Theta_{i}\}_{i=1}^{N}=\{\mu_{i},s_{i},q_{i},\sigma_{i},c_{i}\}_{i=1}^{N}$ are projected to the image plane where their contribution $\alpha_{i}$ is computed from the projected Gaussian (see [77]) and its opacity. For a given pixel, the Gaussians are sorted by depth to compute the pixel color by $\alpha$-blending:
@@
A significant drawback of 3DGS-based approaches is the necessity of a good initialization. State-of-the-art results are only achieved with means $\mu_{i}$ initialized by the sparse depth of Structure-from-Motion [57]. Additionally, heuristics to grow the number of Gaussians to recover under-reconstructed regions are essential. 3DGS initializes new splats based on the gradients of the loss with respect to the camera space means. To address the initialization issues, we generate a prototype of our 3D scene using a text prompt, which we then optimize (Sec. 4.1).
3.2 Diffusion for text-to-3D generation
[…]
Figure 3: Point Cloud for 3D Gaussian Splatting Initalization. We initialize our 3D Gaussian Splatting optimization by leveraging 2D diffusion priors. As shown in a), we first create a reference image $I_{ref}$ from the text prompt $T_{ref}$ at a pre-defined pose $P_{ref}$, and use a monocular depth estimator to convert it to 3D. b) The point cloud is then rendered from neighboring views and expanded to fill in missing regions, as described in Sec. 4.1.
Often, the data distribution is conditional on quantities such as text $T$ and images $I$, so the denoiser takes the form $\epsilon_{\theta}(x_{t},I,T)$ where we absorb the timestep into $x_{t}$. In the conditional case, adherence to conditioning is ensured through Classifier-Free Guidance (CFG), whereby the predicted noise is parameterized as the difference between conditional and unconditional predictions [22, 4]:
@@ 3.2 Diffusion for text-to-3D generation
where $\emptyset$ indicates no conditioning, and the values $S_{I}$ and $S_{T}$ are the guidance weights for image and text, dictating fidelity towards the respective conditions.
Distilling text-to-image diffusion models for text-to-3D generation of object-level data has enjoyed great success since the introduction of Score Distillation Sampling (SDS) [46, 65]. Given a text prompt $T$ and a text-conditioned denoiser $\epsilon_{\theta}(x_{t},T)$, SDS optimizes a 3D model by denoising noised renderings. Given a rendering from a 3D model $x$, we sample a timestep and corresponding $x_{t}$. Considering $\hat{x}=\frac{1}{\alpha_{t}}(x_{t}-\sigma_{t}\epsilon_{\theta}(x_{t},T))$ as the detached one-step prediction of the denoiser, SDS is equivalent to minimizing [76]:
@@
where $w(t)$ is a time-dependent weight over all cameras with respect to the parameters of the 3D representation, and the distribution of $t$ determines the strength of added noise.
@@ 4 Method
We now describe our technique in detail, which broadly consists of three stages: initialization (Fig. 3, Sec. 4.1); inpainting (Fig. 4, Sec. 4.2) with depth distillation (Fig. 4, Sec. 4.3); and finetuning (Fig. 4, Sec. 4.4). Given a text-prompt $T_{\text{ref}}$ and camera poses, we initialize the scene-level 3DGS representation $\{\Theta_{i}\}_{i=1}^{N}$ leveraging 2D diffusion models and monocular depth priors, along with the computed occlusion volume (Sec. 4.1). With this robust initialization, we leverage inpainting diffusion models to fill in disoccluded regions and make a complete 3D scene (Sec. 4.2). In this stage, we also incorporate diffusion depth prior models for better geometry and faster convergence (Sec. 4.3). Finally, we finetune the model with a sharpness filter on sampled images to obtain cohesive 3D samples (Sec. 4.4).
@@ 4.1 Initializing a scene-level 3D representation
Our technique utilizes a text-conditioned optimization procedure to obtain 3DGS scenes, making good initializations essential. A common strategy in this setting is to initialize with a sphere [46, 33] or a low fidelity 3D model provided either by the user [10] or generated via pretrained models [44, 26]. However, such initializations are far from the spatial complexity expected of scenes. Hence, we leverage pretrained 2D priors to create a good initialization, as shown in Fig. 3.
Concretely, we follow an iterative process consisting of image generation, and monocular depth estimation, to build our initial point cloud. We first generate a reference image of the scene $I_{ref}$ from the text prompt $T_{ref}$ with state-of-the-art text-to-image-models. We then use a monocular depth model [29] $\mathcal{D}$ to lift this image to a 3D pointcloud $\mathcal{P}$ from a pre-defined camera pose $P_{ref}$. However, depending on the generated image, the extent of the pointcloud can vary widely. To make the initialization more robust, we outpaint $I_{ref}$ by moving the camera in a small neighborhood of $P_{ref}$ to poses $P_{aux}$, and add points to newly visible areas, similar to [23, 50]. We use an inpainting diffusion model [51] to fill-in these regions and lift this image to 3D using $\mathcal{D}$, while updating $\mathcal{P}$.
Figure 4: Overview of our technique. Our technique first uses a text prompt and an image to build a point cloud (Sec. 4.1) as shown in Fig. 3, which is then completed during the inpainting stage (Sec. 4.2) with an additional depth diffusion prior (Sec. 4.3), and finally a fine-tuning stage (Sec. 4.4) to improve the scene’s coherence.
Given this initial point cloud $\mathcal{P}$, we then precompute the occlusion volume $\mathcal{O}$, which is the set of voxels within the scene bounding box which are occluded by the existing points in $\mathcal{P}$ from $P_{ref}$. We then define the initialization of our 3DGS means as
@@
More details about this process can be found in the supplement.
4.2 Inpainting diffusion models for text-to-3D
@@
Since the pointcloud $\mathcal{P}$ is generated from a a single view and its depth, viewing it from other viewpoints exposes large holes in disoccluded regions. This motivates our formulation of scene generation as an inpainting problem, which we tackle using 2D inpainting diffusion models. Our 2D inpainting diffusion model $\epsilon_{\text{inpaint}}$ [51] takes as input a noisy 3DGS rendering $x_{t}$, the text prompt $T_{\text{ref}}$, an occlusion mask $M_{\text{occl}}$, and the original render $x$. Sampling from this model results in images $\hat{x}$ which fill in the holes in the point cloud rendering while preserving the general structure of the 3D scene, as shown in Fig. 4.
@@
Creating the Inpainting Mask. To compute the conditioning quantities for $\epsilon_{\text{inpaint}}$ from a target camera, we leverage the point cloud $\mathcal{P}$ and the precomputed occlusion volume. Specifically, we render the point cloud and occlusion volume and set all components of $M_{\text{occl}}$ for which the occlusion volume is visible from the target to 00, and $1$ otherwise. Note that this handles cases such as the point cloud occluding itself. We also compute the masked point cloud render $I_{\text{pc}}$ and a mask $M_{\text{val}}$ which is set to 1 for those pixels with corresponding geometry and 0 otherwise.
@@
Computing the Inpainting Loss. We use a 2D inpainting diffusion model which operates in latent space to construct an inpainting loss that provides 3D guidance. The latent diffusion model is defined by its encoder $\mathcal{E}$, decoder $\mathcal{D}$, and the denoising diffusion model $\epsilon_{\text{inpaint}}$. We render an image $x$ with the initialized 3DGS model, and encode it to obtain a latent $z$, where $z=\mathcal{E}(x)$. We then add noise to this latent, yielding $z_{t}$, corresponding to a randomly sampled timestep $t$ from the diffusion model’s noise schedule. Using these quantities, we take multiple DDIM [59] steps from $z_{t}$ to compute a clean latent $\hat{z}$ corresponding to the inpainted image.
Figure 5: Progression of 3D Model after each stage. In this figure, we show how the 3D model changes after each stage in our pipeline. As shown in a) Stage 1 (Sec. 4.1 creates a point cloud with many empty regions. In b), we show the subsequent inpainted model from Stage 2 (Sec. 4.2). Finally, the fine-tuning stage (Sec. 4.4) refines b) to produce the final model, with greater cohesion and sharper detail.
@@
We define our inpainting loss in both latent space and image space, by additionally decoding the predicted latent to obtain $\hat{x}=\mathcal{D}(\hat{z})$. We compute the L2 loss between the latents of the render and sample, as well as an L2 and LPIPS perceptual [74] loss between the rendered image and the decoded sample. To prevent edits outside of the inpainted region, we also add an anchor loss on the unmasked region of $x$, as the L2 difference between $x$ and original point cloud render $I_{pc}$. Our final inpainting loss is
$L$ =$\vphantom{}{}_{\text{inpaint}}=\lambda_{\text{latent}}||z-\hat{z}||_{2}^{2}+\lambda_{\text{image}}||x-\hat{x}||_{2}^{2}$ (5)
$+\lambda_{\text{lpips}}\text{LPIPS}(x,\hat{x})+\lambda_{\text{anchor}}||M_{\text{pc}}(x-I_{\text{pc}})||_{2}^{2}$
with $\lambda$ weighting the different terms and $M_{\text{pc}}=(M_{\text{val}}~{}\&~{}M_{\text{occl}})$ is the bitwise and operation between computed masks. In practice, we also dilate $M_{\text{occl}}$ to improve cohesion at mask boundaries. Further, we find it essential to mask the latent-space L2 loss, to prevent unwanted gradients outside the masked region.
4.3 Depth diffusion for text-to-3D
Recently, image-conditional depth diffusion models have shown state-of-the-art results [29] for relative depth estimation by finetuning a large-scale pre-trained RGB prior [51] on depth datasets. In this section, we show how to distill the knowledge from such methods for text-to-3D synthesis. We consider a depth diffusion model $\epsilon_{\text{depth}}$ which is conditioned by images and text.

### v2

@@ Abstract
We introduce RealmDreamer, a technique for generating forward-facing 3D scenes from text descriptions. Our method optimizes a 3D Gaussian Splatting representation to match complex text prompts using pretrained diffusion models. Our key insight is to leverage 2D inpainting diffusion models conditioned on an initial scene estimate to provide low variance supervision for unknown regions during 3D distillation. In conjunction, we imbue high-fidelity geometry with geometric distillation from a depth diffusion model, conditioned on samples from the inpainting model. We find that the initialization of the optimization is crucial, and provide a principled methodology for doing so. Notably, our technique doesn’t require video or multi-view data and can synthesize various high-quality 3D scenes in different styles with complex layouts. Further, the generality of our method allows 3D synthesis from a single image. As measured by a comprehensive user study, our method outperforms all existing approaches, preferred by 88-95%.
Figure 1: A scene created by our method on the left compared to baseline ProlificDreamer [64] on the right. RealmDreamer generates 3D scenes from text prompts (as above), achieving state-of-the-art results with parallax, detailed appearance, and realistic geometry.
@@ 1 Introduction
Figure 2: Our method, compared to state-of-the-art ProlificDreamer [64] and concurrent work LucidDreamer [14], shows significant improvements. ProlificDreamer yields unsatisfactory geometry and oversaturated renders. LucidDreamer, receiving the same input as our method and an updated depth model [30], displays degeneracy in disoccluded regions, such as the right side of the bed. In contrast, our approach produces visually appealing 3D scenes with realistic geometry.
Text-based 3D scene generation has the potential to revolutionize 3D content creation, with broad applications in virtual reality, game development, and even robotic simulation. However, unlike text-based 2D generative models, 3D data is scarce and lacks diversity, which greatly limits the development of generative 3D techniques. Ideally, one can mitigate this by leveraging rich 2D priors for 3D generation instead. Indeed, object-generation techniques such as DreamFusion [44] and ProlificDreamer [64] do just this, by distilling 2D diffusion priors into a 3D representation, with the latter even demonstrating early abilities to generate scenes. Unfortunately, such distillation approaches can often have saturated results, poor geometry, and lack detail, which become very apparent in the more challenging setting of scene generation (Fig. 2). This leaves the question: How to design a distillation technique for high-quality 3D scene generation from pretrained 2D priors?
A common observation from distillation based object-generation techniques is that greater 3D consistency in 2D diffusion models results in higher-quality distillation, as they provide lower-variance supervision during optimization. As a result, many methods use 2D diffusion models fine-tuned on 3D data [16], such as for novel-view synthesis [37, 21, 45]. Equivalent 3D scene datasets are scarce however, which limits the generalization of such techniques to scenes. Alternatively, ProlificDreamer  [64] fine-tuned a diffusion model during distillation to be more 3D consistent, producing more highly-detailed textures than before. In this work, we introduce a technique to achieve these strengths without requiring 3D training data or fine-tuning existing 2D diffusion models.
We introduce RealmDreamer, a technique for high-fidelity generation of 3D scenes from text prompts (Fig. 1). Our key insight is that we can obtain a 3D scene-aware diffusion model for free, by simply re-appropriating 2D inpainting diffusion models. Typically, 2D inpainting models condition on a partial image to fill in the rest. Instead, we demonstrate that such models can also condition on a 3D scene and fill in unknown regions for novel view synthesis through our proposed inpainting distillation process. As a result, we obtain high-quality 3D scenes with considerably improved detail and appearance over prior distillation techniques. Further, we propose a simple initialization strategy that provides a 3D scene to use as conditioning for this distillation and serves as an initial point cloud for the 3DGS model. We evaluate our technique on several quantitative metrics and obtain significantly higher quality results than prior work, as notably shown by a user study where we are preferred over state-of-the-art ProlificDreamer [64] by 95.5%. Concretely, our contributions are the following:
@@
    An occlusion-aware scene initialization for 3DGS, essential for obtaining high-quality scenes (Sec. 4.1).
@@
    A framework for distillation from 2D inpainting diffusion models which conditions on the existing scene, providing lower variance supervision (Sec. 4.2).
@@
    A method for geometry distillation from diffusion-based depth estimators for higher-fidelity geometry. (Sec. 4.3).
@@
    State-of-the-art results in text-based generation of 3D scenes, as confirmed by several quantitative metrics and a user study (see Fig. 6, Tab. 1, Tab. 2).
@@ 2 Related Work
Figure 3: Overview of our technique. Our technique first uses a text prompt and an image to build a point cloud (Sec. 4.1), which is then completed during the inpainting stage (Sec. 4.2) with an additional depth diffusion prior (Sec. 4.3), and finally a refinement stage (Sec. 4.4) to improve the scene’s coherence.
@@
Text-to-3D. The first methods for text-to-3D generation were based on retrieval from large databases of 3D assets [15, 9, 10]. Subsequently, learning-based methods have dominated  [11, 1, 36]. However, due to the dearth of diverse paired text and 3D data, many recent methods leverage 2D priors, such as CLIP [27, 53] or text-to-image diffusion models  [63, 44, 64, 67, 13, 33]. These distill knowledge from 2D priors into a 3D representation, through variations on Dreamfusion’s score distillation sampling (SDS) [44]. However, these techniques have primarily been limited to object synthesis. In contrast, there are iterative techniques that incrementally build 3D scenes  [26, 14] or 3D-consistent perpetual views [18], but can struggle with high parallax. Our proposed technique builds on strengths from distillation and iterative techniques to produce large scale 3D scenes with high parallax using pretrained 2D priors.
View Synthesis with Diffusion and 3D inpainting. Motivated by the success of SDS, several techniques generate 3D objects from a single image by leveraging image-guided diffusion models to generate novel views and distill to 3D [70, 17]. When trained on larger datasets [16], with better conditioning architectures, these approaches [56, 38, 54, 37, 55] can produce higher quality novel view renders with sharper texture. Some methods also condition denoising directly on renderings from 3D consistent models [21, 8] for view synthesis in a multi-view consistent manner. Unfortunately, most techniques rely on object-level data, limiting their use for text-based scene synthesis. 3D inpainting techniques [42, 41] also leverage image-guided diffusion models to remove small objects in scenes. Other works focus on training custom inpainting models for indoor scenes [32] or objects [28] to generate novel views. In contrast to these, we leverage pre-trained text-guided inpainting priors and focus on generating large missing regions of diverse scenes with our novel inpainting distillation loss.
Concurrent work. In the rapidly evolving text-to-3D field, we focus on the most relevant concurrent works, highlighting our key differences. LucidDreamer [14] and Text2NeRF [68] uses an iterative approach similar to PixelSynth [49] and Text2Room [26] to generate 3D scenes but displays limited parallax. Considering LucidDreamer as the most relevant concurrent baseline, we compare it in the fairest setting possible, by using newer depth estimators [30, 65], and surpass it by $88.5\%$ in our user study. Most recently, in follow-up work, CAT3D [20], utilizes a diffusion model finetuned on multiview datasets to generate multiple views from a single image. In contrast, our entire pipeline does not use multiview images.
@@ 3 Preliminaries
The splats $\{\Theta_{i}\}_{i=1}^{N}=\{\mu_{i},s_{i},q_{i},\sigma_{i},c_{i}\}_{i=1}^{N}$ are projected to the image plane where their contribution $\alpha_{i}$ is computed from the projected Gaussian (see [72]) and $\sigma_{i}$. A pixel’s color is obtained by $\alpha$-blending Gaussians sorted by depth:
@@
A significant drawback of 3DGS-based approaches is the necessity of a good initialization. State-of-the-art results are only achieved with means $\mu_{i}$ initialized by the sparse depth of Structure-from-Motion [57], which is not applicable for scene generation. To address this challenge, we generate a prototype of our 3D scene using a text prompt, which we then optimize (Sec. 4.1).
Figure 4: Progression of 3D Model after each stage. We show how the 3D model changes after each stage in our pipeline. As shown in a) Stage 1 (Sec. 4.1) creates a point cloud with many empty regions. In b), we show the subsequent inpainted model from Stage 2 (Sec. 4.2). Finally, the fine-tuning stage (Sec. 4.4) refines b) to produce the final model, with greater cohesion and sharper detail.
3.2 Conditional Diffusion Models
Diffusion models [58, 25, 29, 59, 61, 60] are generative models which learn to map noise $x_{T}\sim\mathcal{N}(0,I)$ to data by iteratively denoising a set of latents $x_{t}$ corresponding to decreasing noise levels $t$ using non-deterministic DDPM [25] or deterministic DDIM sampling [59], among others [29, 61, 60].
Given $t$, a diffusion model $\epsilon_{\theta}$ is trained to predict the noise $\epsilon$ added to the image such that we obtain $\epsilon_{\theta}(x_{t},t)$, which approximates the direction to a higher probability density. Often, the data distribution is conditional on quantities such as text $T$ and images $I$, so the denoiser takes the form $\epsilon_{\theta}(x_{t},I,T)$. In the conditional case, classifier-free guidance is often used to obtain the predicted noise [24, 6]:
@@ 3.2 Conditional Diffusion Models
where $\emptyset$ indicates no conditioning, and the values $S_{I}$ and $S_{T}$ are the guidance weights for image and text, dictating fidelity towards the respective conditions. In the case of latent diffusion models like Stable Diffusion [50], denoising happens in a compressed latent space by encoding and decoding images with an encoder $\mathcal{E}$ and decoder $\mathsf{D}$.
Score Distillation Sampling. Distilling text-to-image diffusion models for text-to-3D generation of object-level data has enjoyed great success since the introduction of Score Distillation Sampling (SDS) [44, 63]. Given a text prompt $T$ and a text-conditioned denoiser $\epsilon_{\theta}(x_{t},T)$, SDS optimizes a 3D model by denoising noised renderings. Given a rendering from a 3D model $x$, we sample a timestep and corresponding $x_{t}$. Considering $\hat{x}=\frac{1}{\alpha_{t}}(x_{t}-\sigma_{t}\epsilon_{\theta}(x_{t},T))$ as the detached one-step prediction of the denoiser, SDS is equivalent to minimizing [71]:
@@
where $w(t)$ is a time-dependent weight over all cameras with respect to the parameters of the 3D representation, and the distribution of $t$ determines the strength of added noise. In this work, we use a variation of SDS to distill from pretrained-inpainting models (Sec. 4.2)
@@ 4 Method
We now describe our technique in detail, which broadly consists of three stages: initialization (left of Fig. 3, Sec. 4.1); inpainting (middle of Fig. 3, Sec. 4.2) with depth distillation (middle of Fig. 3, Sec. 4.3); and finetuning (right of Fig. 3, Sec. 4.4). Given a text-prompt $T_{\text{ref}}$ and camera poses, we initialize the scene-level 3DGS representation $\{\Theta_{i}\}_{i=1}^{N}$ leveraging 2D diffusion models and monocular depth priors, along with the computed occlusion volume (Sec. 4.1). With this robust initialization, we use 2D inpainting models to predict novel views, distilling to 3D to create a complete 3D scene (Sec. 4.2). In this stage, we also incorporate depth distillation for higher-quality geometry (Sec. 4.3). Finally, we refine the model with a sharpness filter on sampled images to obtain high-quality 3D samples (Sec. 4.4). The result from these stages are shown in Fig. 4.
Figure 5: Qualitative Results. In the left column, we show the input prompt for our technique. In the next two columns, we show the renderings from our 3D model from different viewpoints. In the fourth column, we show the level of agreement between rendering and geometry by a split view of the rendering and depth. Finally, in the last column, we show the depth map.
@@ 4.1 Initializing a Scene-level 3D Representation
Our technique utilizes 3DGS for text-conditioned optimization, making a good initialization essential. A common strategy in this setting is to initialize with a sphere [44, 34] but the density of a scene is more complex and distributed. Hence, we leverage pretrained 2D priors to synthesize a robust initialization (left of Fig. 3).
Concretely, we first generate a reference image of the scene $I_{ref}$ from the text prompt $T_{ref}$ with a state-of-the-art text-to-image-model. We then employ a monocular depth model [30] $\mathcal{D}$ to lift this image to a pointcloud $\mathcal{P}$ from corresponding camera pose $P_{ref}$. Depending on the generated image, the extent of the pointcloud can vary widely. To make the initialization more robust, we outpaint $I_{ref}$ by moving the camera left and right of $P_{ref}$ to poses $P_{aux}$. We use an inpainting diffusion model [50] to fill in the unseen regions which are lifted to 3D using $\mathcal{D}$. The union of all generated points thus becomes $\mathcal{P}$.
Determining Incomplete Regions. Given the initial point cloud $\mathcal{P}$, we then precompute the undetermined 3D region, or the occlusion volume $\mathcal{O}$, which is the set of voxel centers within the scene’s occupancy grid which are occluded by the existing points in $\mathcal{P}$ from $P_{ref}$. We use $\mathcal{O}$ when computing inpainting masks later and define the initialization of our 3DGS means as
@@
More details can be found in the supplementary.
4.2 Inpainting Diffusion for 3D-Conditioned Distillation
@@
Since our initialization is generated from sparse poses, viewing it from novel viewpoints exposes large holes in disoccluded regions (Fig. 4). We resolve this with a novel inpainting distillation technique, that conditions a 2D inpainting diffusion model $\epsilon_{\text{inpaint}}$ [50] on the existing scene to complete missing regions. The model takes as input a noisy rendering $x_{t}$ of $\{\Theta_{i}\}_{i=1}^{N}$, and is conditioned by the text prompt $T_{\text{ref}}$, an occlusion mask $M_{\text{occl}}$, and the point cloud render $I_{\text{pc}}$. Sampling from this model results in novel views $\hat{x}$ which plausibly fill in the holes in the renderings while preserving the structure of the 3D scene (Fig. 3).
@@
Conditioning the inpainting model. To compute the conditioning mask $M_{\text{occl}}$ for $\epsilon_{\text{inpaint}}$, we render the point cloud $\mathcal{P}$ and the precomputed occlusion volume $\mathcal{O}$. We set all components of $M_{\text{occl}}$ for which the occlusion volume is visible from the target to 00, and $1$ otherwise. Note that this handles cases such as the point cloud occluding itself (see the supplement for a visualization).
@@
Computing the inpainting loss. Our 2D inpainting diffusion model $\epsilon_{\text{inpaint}}$ [50] operates in latent space, thus additionally parametrized by its encoder $\mathcal{E}$ and decoder $\mathsf{D}$. We render an image $x$ with the initialized 3DGS model, and encode it to obtain a latent $z$, where $z=\mathcal{E}(x)$. We then add noise to this latent, yielding $z_{t}$, corresponding to a randomly sampled timestep $t$ from the diffusion model’s noise schedule. Using these quantities, we take multiple DDIM [59] steps from $z_{t}$ to compute a clean latent $\hat{z}$ corresponding to the inpainted image.
Figure 6: Qualitative Comparisons. Our technique shows superior quality in appearance and geometry than all baselines. Please see the supplementary for more comparisons. Prompt: “A boy sitting in a boat in the middle of the ocean, under the milkyway, anime style”.
@@
We define our inpainting loss in both latent space and image space, by additionally decoding the predicted latent to obtain $\hat{x}=\mathsf{D}(\hat{z})$. We compute the L2 loss between the latents of the render and sample, as well as an L2 and LPIPS perceptual [69] loss between the rendered image and the decoded sample. To prevent edits outside of the inpainted region, we also add an anchor loss on the unmasked region of $x$, as the L2 difference between $x$ and original point cloud render $I_{pc}$. Our final inpainting loss is
$L$ =${}_{\text{inpaint}}=\lambda_{\text{latent}}||z-\hat{z}||_{2}^{2}+\lambda_{\text{image}}||x-\hat{x}||_{2}^{2}$ (5)
$+\lambda_{\text{lpips}}\text{LPIPS}(x,\hat{x})+\lambda_{\text{anchor}}||M_{\text{occl}}(x-I_{\text{pc}})||_{2}^{2}$
with $\lambda$ weighting the different terms. We discuss the similarity of this loss with SDS in the supplemetary.
Discussion. In contrast to existing iterative methods which utilize inpainting (such as Text2Room and LucidDreamer), our framework does not iteratively construct a scene with inpainting. In practice, sampling from inpainting models often produces artifacts (such as due to out-of-distribution masks), which iterative approaches can amplify when generating from new poses. In contrast, due to scene-conditioned multiview optimization, we obtain cohesive 3D scenes and do not progressively accumulate errors. Moreover, in contrast to DreamFusion and ProlificDreamer, our method utilizes a scene-conditional diffusion model, providing lower variance updates for effective optimization (see row 2 of Fig. 7). This avoids the high-saturation and blurry results that are typically found (Fig. 6).
4.3 Depth Diffusion for Geometry Distillation
To improve the quality of generated geometry, we incorporate a pretrained geometric prior to avoid degenerate solutions. Here, we leverage monocular depth diffusion models and propose an additional depth distillation loss (middle of Fig. 3). Crucially, we integrate this with our inpainting distillation by conditioning the depth model $\epsilon_{\text{depth}}$ on the aforementioned samples $\hat{x}$ from $\epsilon_{\text{inpaint}}$.